### v1

@@ III Helmholtz free energy and Gibbs state
Considering a system at fixed temperature $T>0$ (through coupling to a heat bath) means that we cannot rely on a variational principle for the internal energy $E_{v}(\Gamma)=\operatorname{Tr}\{\Gamma H_{v}\}$ any more [19]. Instead, we have to consider the Helmholtz free energy
@@
where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The new variational principle then yields the Helmholtz functional
$\Omega^{\beta}(v)=\inf_{\Gamma}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma}\operatorname{Tr}\bigl{\{}\Gamma\bigl{(}H_{v}+\beta^{-1}\log(\Gamma)\bigr{)}\bigr{\}}$ (12)
and can been equivalentely seen as the Legendre transform of $E_{v}$ (in terms of entropy) to $\Omega^{\beta}(v)$ (in terms of temperature). Here the infimum is over all density matrices, i.e. over all (possibly infinite) convex sums $\Gamma=\sum_{k}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}$, $\lambda_{k}\geq 0$, $\sum_{k}\lambda_{k}=1$, with $\Psi_{k}\in\mathscr{W}$ all orthogonal. Making the Helmholtz functional stationary with respect to small perturbation $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
@@
The unique minimizer of the Helmholtz functional is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$ where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function. The state $\Gamma_{v}$ is called the Gibbs state and it does only exist if $Z(v)$ is finite. Moreover, for each $v\in H^{-1}(\mathbb{T})$ the Gibbs state exists and the Helmholtz free energy is given by
$\Omega^{\beta}(v)=-\beta^{-1}\log\bigl{(}Z(v)\bigr{)}.$ (13)
The proof of Proposition 3 can be found in Appendix A. The reason for $Z(v)$ to be finite is that the particles are trapped in a finite volume. The defining parameters of canonical (quantum) thermodynamics are $(N,V,T)$ and all of these parameters need to be finite. For a counterexample [16, Sec. 2.3] consider the hydrogen atom in $\mathbb{R}^{3}$. The energies are $1/(2n^{2})$ with a degeneracy of $n^{2}$ and with $n\in\mathbb{N}$. Then the partition function is
@@
Proposition 3 implies that the density of the Gibbs state is an infinite convex combination of elements in $\mathscr{I}$. Hence we cannot directly conclude that its density is contained in $\mathscr{I}$, but we will show this statement with the following proposition.
@@
The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}$.
The proposition is proven in Appendix B. It tells us that $v$-representable densities at elevated temperatures are contained in $\mathscr{I}$. Hence we can split the minimization of the Helmholtz functional $\Omega^{\beta}_{v}$ as in the zero temperature case into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density,
@@
$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\Gamma\mapsto\rho}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (15b)
The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$. Note that because of the existence of the Gibbs state, the infimum in the variation over the densities is attained whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
As for zero temperature ([6; 7]), we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions. Due to the concavity of the entropy [16, Th. 12] and linearity of $\Gamma\to\rho$ the following proposition follows.
@@
    The Helmholtz functional $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex.
@@
    The universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
IV Relation between $v$-representability and subdifferentials
The relation between $v$-representability and non-empty subdifferential was presented in the literature [20; 21; 22]. This relation, together with some additional results from convex analysis, are important ingredients to show our main result Theorem 1. Therefore, we will repeat the relation and the needed results in this section.
@@ IV Relation between $v$-representability and subdifferentials
Note, that if the inequality $f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle$ cannot be satisfied for any $u\in\mathscr{X}^{*}$ and all $\rho^{\prime}\in\mathscr{X}$ then the subdifferential is empty. An element of $\partial f(\rho)$ is called a subgradient. Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [23, Prop. 2.40]. In this case, this element is equal to the derivative. If the general convex function is replaced by the universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
@@
In particular, $\rho$ is a minimizer for the variation in Eq. (15a). Note that this is only a necessary condition for $v$-representability and not a sufficient one. The reason is that for $v$-representability of a density $\rho$ we additionally require that the variation over $\Gamma\mapsto\rho$ in Eq. (15b) has a minimum too, otherwise there is no Gibbs state behind the density. In this section, we show that the subdifferential of $F^{\beta}_{\mathrm{DM}}(\rho)$ at any density $\rho\in\mathscr{X}_{>0}$ is non-empty. In the next section we then show that these densities indeed come from Gibbs states. The two result together give then $v$-representability for all densities in $\mathscr{X}_{>0}$.
To show the existence of a subgradient for a density $\rho\in\mathscr{X}_{>0}$ we proceed as in Sutter et al. [7]. The only result that significantly changes compared to this reference is the local boundedness of the universal functional $F_{\mathrm{DM}}^{\beta}$.
@@
$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0)$ (19)
which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta S(\Gamma)$ is negative. Hence we get the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\Gamma\to\rho}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2}$ (20)
where we used a previous result [7, Lem. 10] for the second inequality. ∎
@@
For any $\rho\in\mathscr{X}_{>0}$ the universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
The proof is the same as in Sutter et al. [7, Th. 15].
@@ V Existence of the Minimum
As discussed in the previous section, we do not only need a non-empty subdifferential of $F_{\mathrm{DM}}^{\beta}(\rho)$ to get $v$-representability but also that the minimum in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ (Eq. (15b)) exists. This minimizer must then be the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$. In the following we show this for all $\rho\in\mathscr{X}_{>0}$. To do so, we make use of the relative entropy which is defined as
$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}$ (21)
for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$. An important property of the relative entropy is lower semi-continuity.
@@
The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology. (for non-negative trace class operators).
The next lemma relates the Helmholtz functional and the relative entropy. Note that originally we only defined the Helmholtz functional $\Omega^{\beta}_{v}(\Gamma)$ for (normalized) density matrices. However, in the same manner it can be defined for any non-negative trace class operator. Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and it is given by $\mathrm{e}^{-\beta H_{v}}/Z(v)$.
@@
Since the left hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (norm is decreasing under weak-* limits), we get that the Helmholtz functional together with a term proportional to $\operatorname{Tr}\{\cdot\}$ is lower semi-continuous with respect to the weak-* topology. In particular, we get the following lemma.
@@
For any $\rho\in\mathscr{X}_{>0}$ there is a (normalized) density matrix $\Gamma\mapsto\rho$ such that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log\Gamma)\}$. This shows the existence of a minimizer in Eq. (15b).
@@
For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exists by Theorem 7. By definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ there is a sequence of density matrices $\Gamma_{n}$ each yielding the density $\rho$ and such that $\operatorname{Tr}\{\Gamma_{n}(H_{0}+\beta^{-1}\log\Gamma_{n})\}\to F^{\beta}_{\mathrm{DM}}(\rho)$. Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [18, Th. IV.21] that there is a subsequence (in the following also denoted by $\Gamma_{n}$) and a trace class operator $\Gamma_{0}$ such that $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$. Moreover, we know that $\Gamma_{0}$ is non-negative and that $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$. To prove the theorem, we show that $\Gamma_{0}$ is a minimizer. That means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$ and that it yields the density $\rho$. We start by showing $\operatorname{Tr}\{\Gamma_{0}\}=1$.
For simplicity, we call the minimum of the Helmholtz functional $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$. Because $v\in\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
@@
$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\}$
where we have used the fact that the function $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous due to Lemma 10. For the condition of the lemma to be fulfilled we set $\xi=(\beta\Omega^{\beta})^{-1}+1$. This expression is well-defined, since we can choose a representative of $v]\in\mathscr{X}^{*}$ for which $\Omega^{\beta}_{v}(\Gamma_{v})$ is non-zero. So the condition of Lemma 10 is satisfied. Let $c$ be the constant such that $\operatorname{Tr}\{c\Gamma_{0}\}=1$. (Note that the constant $c$ only exist if $\Gamma_{0}\neq 0$. If $\Gamma_{0}=0$ then the right hand-side of Eq. (23) is 0 where the left hand-side is $\beta^{-1}$. Hence $\Gamma_{0}=0$ is not possible.) Since $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$ it holds $c\geq 1$. Then we can compute
@@
$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c$
where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq. (23). This inequality can be transformed to
@@
It is clear that we have equality if $c$ is equal to $1$. The function $f(c)=c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}$ defined on $1,\infty)$ has $f(1)=0$ and the derivative
@@
Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\to\rho$. For this we use the same idea as Lieb [6] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$ where $M_{f}$ denotes pointwise multiplication by $f$. This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\to\rho$. To do this, let $|\sum_{j}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of $\Gamma_{0}$. Since $\operatorname{Tr}\{\Gamma_{0}\}=1$ there is for any $\epsilon>0$ an $N_{\epsilon}\in\mathbb{N}$ such that
@@
The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ in trace class, there exist $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$. It follows that for $n>N_{P}$ we have $\operatorname{Tr}\{\Gamma_{n}P\}>\operatorname{Tr}\{\Gamma_{0}P\}-\epsilon>1-2\epsilon$ and thus $\operatorname{Tr}\{\Gamma_{n}(1-P)\}<2\epsilon$. Then we get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$. We can write

### v2

@@ III Helmholtz free energy and Gibbs state
Considering a system at fixed temperature $T>0$ (through coupling to a heat bath) means that we cannot rely on a variational principle for the internal energy $E_{v}(\Gamma)=\operatorname{Tr}\{\Gamma H_{v}\}$ any more [3]. Instead, we have to consider the Helmholtz free energy for arbitrary density matrices,
@@
where $S(\Gamma)=-\operatorname{Tr}\{\Gamma\log(\Gamma)\}$ is the entropy and $\beta=T^{-1}$ is the inverse temperature. The variational principle then yields the Helmholtz functional,
$\Omega^{\beta}(v)=\inf_{\Gamma\in\mathscr{D}}\Omega^{\beta}_{v}(\Gamma)=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\bigl\{\Gamma\bigl(H_{v}+\beta^{-1}\log(\Gamma)\bigr)\bigr\}.$ (18)
Here the infimum is taken over the set of all density matrices,
$\mathscr{D}=\left\{\Gamma=\sum_{k=1}^{\infty}\lambda_{k}\langle\Psi_{k},\cdot\rangle\Psi_{k}\;\middle|\;\lambda_{k}\geq 0,\operatorname{Tr}\{\Gamma\}=1,\Psi_{k}\in\mathscr{W}\;\text{all orthogonal},S(\Gamma)<\infty\right\}.$ (19)
Making the Helmholtz free energy stationary with respect to small perturbations $\delta\Gamma$ with $\operatorname{Tr}\{\delta\Gamma\}=0$ gives the minimizer.
@@
The unique minimizer of the Helmholtz free energy is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)\in\mathscr{D}$, where $Z(v)=\operatorname{Tr}\{\mathrm{e}^{-\beta H_{v}}\}$ is called the partition function. The state $\Gamma_{v}$ is called the Gibbs state and only exists if $Z(v)$ is finite. Moreover, for each $v\in H^{-1}(\mathbb{T})$ the Gibbs state exists and the Helmholtz free energy is given by
$\Omega_{v}^{\beta}(\Gamma_{v})=\Omega^{\beta}(v)=-\beta^{-1}\log\bigl(Z(v)\bigr).$ (20)
Remark 1.
For an example of a physical system with infinite partition function simply consider the hydrogen atom in the infinite-volume space $\mathbb{R}^{3}$ [17, Sec. 2.3]. The eigenenergies are $1/(2n^{2})$ with multiplicity $^{2}, nn\in\mathbb{N}$. Then the partition function is
@@
The proof of Proposition 3 can be found in Appendix A. Proposition 3 implies that the density of the Gibbs state is an infinite convex combination of elements in $\mathscr{I}$. Hence we cannot directly conclude that its density is contained in $\mathscr{I}$, but we will show this statement with the following proposition.
@@
The density of the Gibbs state $\Gamma_{v}$ is contained in $\mathscr{I}$ for all $v\in H^{-1}(\mathbb{T})$.
The proposition is proven in Appendix B. It tells us that $v$-representable densities at elevated temperatures are always contained in $\mathscr{I}$. Hence we can split the minimization of the Helmholtz free energy $\Omega^{\beta}_{v}$ into first a variation over densities in $\mathscr{I}$ followed by a minimization over all density matrices which yield such a density, just like in the zero temperature case,
@@
$F^{\beta}_{\mathrm{DM}}(\rho)=\inf_{\begin{subarray}{c}\Gamma\in\mathscr{D}\\ \Gamma\mapsto\rho\end{subarray}}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}.$ (22b)
The functional $F_{\mathrm{DM}}^{\beta}(\rho)$ is called the thermal universal functional and we can extend it to $\mathscr{X}$ by setting $F_{\mathrm{DM}}^{\beta}(\rho)=\infty$ for all $\rho\in\mathscr{X}\setminus\mathscr{I}$. Note that because of the existence of the Gibbs state and Propositions 3 and 4, the infimum in the variation over the densities is always attained, whereas for the second infimum (in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$) we only know about the existence of a minimizer for $v$-representable densities which are characterized in Theorem 1.
Like in the zero-temperature case [28; 38], we would like to have a convex formulation of thermal density-functional theory to exploit the properties of convex functions. Due to the concavity of the entropy [17, Th. 12] and linearity of $\Gamma\to\rho$ the following proposition follows.
@@
    The Helmholtz free energy $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)$ is convex on $\mathscr{D}$.
@@
    The thermal universal functional $\rho\mapsto F^{\beta}_{\mathrm{DM}}(\rho)$ is convex on $\mathscr{X}$.
IV Relation between $v$-representability and non-empty subdifferential
The relation between $v$-representability and non-empty subdifferential is usually discussed in mathematical presentations of DFT [12; 32; 6]. This relation, together with some additional results from convex analysis, are important ingredients to show our main Theorem 1. Therefore, we will repeat the relation and the needed results in this section.
@@ IV Relation between $v$-representability and non-empty subdifferential
Note, that if the inequality $f(\rho)-f(\rho^{\prime})\leq\langle u,\rho-\rho^{\prime}\rangle$ cannot be satisfied for any $u\in\mathscr{X}^{*}$ and all $\rho^{\prime}\in\mathscr{X}$ then the subdifferential is empty. An element of $\partial f(\rho)$ is called a subgradient. Moreover, the function $f$ is Gâteaux differentiable at a point $\rho$ if the function $f$ is continuous at $\rho$ and if the subdifferential $\partial f(\rho)$ contains only a single element [2, Prop. 2.40]. In this case, this element is equal to the derivative. If the general convex function is replaced by the thermal universal functional (which is a proper convex functional by Proposition 5), we get for an element $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ and for all $\rho^{\prime}\in\mathscr{X}$ the following inequality
@@
In particular, $\rho$ is a minimizer for the variation in Eq. (22a). Note that this is only a necessary condition for $v$-representability and not a sufficient one. The reason is that for $v$-representability of a density $\rho$ we additionally require that the variation over $\Gamma\mapsto\rho$ in Eq. (22b) has a minimum too, otherwise there is no Gibbs state behind the density. In this section, we show that the subdifferential of $F^{\beta}_{\mathrm{DM}}(\rho)$ at any density $\rho\in\mathscr{X}_{>0}$ is non-empty. In the next section we then show that these densities indeed come from Gibbs states. Only the two result together then give $v$-representability for all densities in $\mathscr{X}_{>0}$.
To show the existence of a subgradient for a density $\rho\in\mathscr{X}_{>0}$ we proceed as in Sutter et al. [38]. The only result that significantly changes compared to our previous work is the local boundedness of the thermal universal functional $F_{\mathrm{DM}}^{\beta}$.
@@
$F_{\mathrm{DM}}^{\beta}(\rho)\geq\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{0}+\beta^{-1}\log(\Gamma))\}=\Omega^{\beta}(0),$ (26)
which is greater than -$\infty$ by Proposition 3. For the upper bound we note that the entropic term $\beta^{-1}S(\Gamma)$ is negative. Hence we get the following inequality
$F_{\mathrm{DM}}^{\beta}(\rho)\leq\inf_{\begin{subarray}{c}\Gamma\in\mathscr{D}\\ \Gamma\mapsto\rho\end{subarray}}\operatorname{Tr}\{\Gamma H_{0}\}\leq C_{2}+C_{3}\|\nabla\sqrt{\rho}\|_{L^{2}}^{2},$ (27)
where we used a previous result [38, Lem. 10] for the second inequality. ∎
@@
At any $\rho\in\mathscr{X}_{>0}$ the thermal universal functional $F_{\mathrm{DM}}^{\beta}$ is continuous and the subdifferential $\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$ is non-empty.
The proof is the same as in Sutter et al. [38, Th. 15]. The result relies critically on the embedding $H^{1}\hookrightarrow L^{\infty}$ from Theorem 2(iii) that can be used to show that $\mathscr{X}_{>0}$ is an open set [38, Lem. 14]. This embedding only holds true in one dimension and one further needs the torus setting with a non-vanishing minimum of $\rho$ in order to show that $F_{\mathrm{DM}}^{\beta}$ is bounded above. Both results combined then yield the desired result by using powerful theorems from convex analysis [2, Th. 2.14 and Prop. 2.36]. We note that the same result can be established also on the open interval with Neumann boundary conditions [8]. We can now also understand the origin of the distributional potentials. We have $\mathscr{X}_{>0}\subset\mathscr{X}\subset H^{1}(\mathbb{T})$, so the density domain is limited to a subset of the Sobolev space $H^{1}(\mathbb{T})$ with dual space $H^{-1}(\mathbb{T})\subset\mathscr{X}^{*}$ that contains distributions. On the other hand, a potential is always from the subdifferential of the thermal universal functional and the subdifferential itself is a subset of the dual space $\mathscr{X}^{*}$.
@@ V Existence of the Minimum
As discussed in the previous section, we do not only need a non-empty subdifferential of $F_{\mathrm{DM}}^{\beta}(\rho)$ to get $v$-representability but also that the minimum in the definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ (Eq. (22b)) exists. This minimizer is then the Gibbs state $\Gamma_{v}$, where $v$ is just the potential that yields $\rho$. In the following we show this for all $\rho\in\mathscr{X}_{>0}$. To do so, we make use of the relative entropy which for two non-negative trace class operators $A$, $B$ with $\ker B\subset\ker A$ is defined as
$S(A|B)=\operatorname{Tr}\{A(\log A-\log B)+B-A\}.$ (28)
The relative entropy is always non-negative, $S(A|B)\geq 0$ and $S(A|B)=0$ if and only if $A=B$ [29; 14]. An important property of the relative entropy is its lower semi-continuity.
@@
The relative entropy $(A,B)\mapsto S(A|B)$ is jointly lower semi-continuous with respect to the weak-* topology.
The next lemma relates the Helmholtz free energy and the relative entropy. Note that originally we only defined the Helmholtz free energy $\Omega^{\beta}_{v}(\Gamma)$ for $\Gamma\in\mathscr{D}$. However, in the same manner, it can be defined for any non-negative trace-class operator with finite entropy. Recall that $\Gamma_{v}$ is the Gibbs state corresponding to a potential $v$ and is given by $\Gamma_{v}=\mathrm{e}^{-\beta H_{v}}/Z(v)$. Its kernel is trivial and thus $S(\cdot|\Gamma_{v})$ is well-defined.
@@
Note that the left-hand side is well-defined for all non-negative trace-class operators $\Gamma$ and thus Lemma 9 extends the domain of $\Omega_{v}^{\beta}(\cdot)$ to density matrices with infinite entropy. Since the left-hand side in Lemma 9 consists of two weak-* lower semi-continuous functions (the trace norm $\operatorname{Tr}|\Gamma|$ is decreasing under weak-* limits and equal to $\operatorname{Tr}\{\Gamma\}$ since $\Gamma\geq 0$), we get that the Helmholtz free energy together with a term proportional to $\operatorname{Tr}\{\cdot\}$ (with positive factor) is lower semi-continuous with respect to the weak-* topology. In particular, we get the following lemma.
@@
For every $\rho\in\mathscr{X}_{>0}$ and all potentials $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)\subset\mathscr{X}^{*}$, it holds that the corresponding Gibbs state $\Gamma_{v}$ yields the density $\rho$ and is the minimizer in Eq. (22b), i.e., $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{v}(H_{0}+\beta^{-1}\log\Gamma_{v})\}$.
@@
For a density $\rho\in\mathscr{X}_{>0}$ we pick $v\in-\partial F_{\mathrm{DM}}^{\beta}(\rho)$ which we know to exist by Theorem 7. By definition of $F_{\mathrm{DM}}^{\beta}(\rho)$ there is a sequence of density matrices $\Gamma_{n}$ each yielding the density $\rho$ and such that $\operatorname{Tr}\{\Gamma_{n}(H_{0}+\beta^{-1}\log\Gamma_{n})\}\to F^{\beta}_{\mathrm{DM}}(\rho)$. Since all $\Gamma_{n}$ are contained in a ball of radius $1$ we get by the Banach–Alaoglu theorem [35, Th. IV.21] that there is a subsequence (in the following also denoted by $\Gamma_{n}$) and a trace class operator $\Gamma_{0}$ such that $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$. Moreover, we know that $\Gamma_{0}$ is non-negative and that $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$. To prove the theorem, we show that $\Gamma_{0}$ is a minimizer in Eq. (22b) that, together with the discussion after Eq. (24), already shows that $\Gamma_{0}$ is equal to the Gibbs state $\Gamma_{v}\in\mathscr{D}$. This means that $\Gamma_{0}$ is normalized to $1$, that $F_{\mathrm{DM}}^{\beta}(\rho)=\operatorname{Tr}\{\Gamma_{0}(H_{0}+\beta^{-1}\log\Gamma_{0})\}$, and that it yields the density $\rho$. We start by showing $\operatorname{Tr}\{\Gamma_{0}\}=1$.
For simplicity, we call the minimum of the Helmholtz free energy $\Omega^{\beta}=\Omega^{\beta}_{v}(\Gamma_{v})=\inf_{\Gamma\in\mathscr{D}}\operatorname{Tr}\{\Gamma(H_{v}+\beta^{-1}\log\Gamma)\}$. Because $v\in-\partial F^{\beta}_{\mathrm{DM}}(\rho)$ and $\lim_{n\to\infty}\operatorname{Tr}\{\Gamma_{n}\}=1$ we have
@@
$\geq\operatorname{Tr}\{\Gamma_{0}(H_{v}+\beta^{-1}\log\Gamma_{0})\}-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma_{0}\},$
where we have used the fact that the function $\Gamma\mapsto\Omega^{\beta}_{v}(\Gamma)-\xi\Omega^{\beta}\operatorname{Tr}\{\Gamma\}$ is weak-* lower semi-continuous due to Lemma 10. For the condition of the lemma to be fulfilled we set $\xi=(\beta\Omega^{\beta})^{-1}+1$. This expression is well-defined, since we can choose a representative of $v]\in\mathscr{X}^{*}$ for which $\Omega^{\beta}_{v}(\Gamma_{v})$ is non-zero. Let $c$ be the constant such that $\operatorname{Tr}\{c\Gamma_{0}\}=1$. (Note that the constant $c$ only exist if $\Gamma_{0}\neq 0$. If $\Gamma_{0}=0$ then the right-hand side of Eq. (30) is 0 while the left-hand side is $\beta^{-1}$. Hence $\Gamma_{0}=0$ is not possible.) Since $\operatorname{Tr}\{\Gamma_{0}\}\leq 1$ it holds $c\geq 1$. We compute
@@
$\leq c(\Omega^{\beta}-\xi\Omega^{\beta})+\beta^{-1}\log c,$
where we used the variational principle for $\Omega^{\beta}\leq\Omega^{\beta}_{v}(c\Gamma_{0})$, the fact that $\operatorname{Tr}\{c\Gamma_{0}\}=1$, and Eq. (30). This inequality can be transformed into
@@
It is clear that we have equality if $c=1$. The function $f(c)=c\Omega^{\beta}-c\xi\Omega^{\beta}+\beta^{-1}\log c-\Omega^{\beta}+\xi\Omega^{\beta}$ defined on $1,\infty)$ has $f(1)=0$ and the derivative
@@
Now that we know that the weak-* limit $\Gamma_{0}$ has trace norm equal to $1$ we can show that $\Gamma_{0}\mapsto\rho$. For this we use the same idea as Lieb [28] and show that for any $f\in L^{\infty}(\mathbb{T})$ we have $\operatorname{Tr}\{\Gamma_{n}M_{f}\}\to\operatorname{Tr}\{\Gamma_{0}M_{f}\}$, where $M_{f}$ denotes pointwise multiplication by $f$. This means that the densities of $\Gamma_{n}$ converge weakly to the density of $\Gamma_{0}$ (in $L^{1}$) and by the uniqueness of weak limits we get $\Gamma_{0}\mapsto\rho$. To do this, let $|\sum_{j}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$ be the spectral decomposition of $\Gamma_{0}$. Since $\operatorname{Tr}\{\Gamma_{0}\}=1$ there is for any $\epsilon>0$ an $N_{\epsilon}\in\mathbb{N}$ such that
@@
The operator $P$ is compact and since $\Gamma_{n}\overset{\ast}{\rightharpoonup}\Gamma_{0}$ is trace class there exists an $N_{P}$ such that $\lvert\operatorname{Tr}\{\Gamma_{n}P\}-\operatorname{Tr}\{\Gamma_{0}P\}\rvert<\epsilon$ for all $n>N_{P}$. It follows that for $n>N_{P}$ we have $\operatorname{Tr}\{\Gamma_{n}P\}>\operatorname{Tr}\{\Gamma_{0}P\}-\epsilon>1-2\epsilon$ and thus $\operatorname{Tr}\{\Gamma_{n}(1-P)\}<2\epsilon$. We thus get that $\lvert\operatorname{Tr}\{(\Gamma_{n}-\Gamma_{0})(1-P)\}\rvert<2\epsilon$ for large enough $n$ and we can write